if a baby is randomly selected, find the probability that the baby is a female or has black hair?
how am i solve to this? she gives us a chart

Brown Blonde Black Auburn None Totals
Male 5 24 28 0 8 65
Female 9 18 16 3 5 51
Totals 14 42 44 3 13 116

Answer by ewatrrr(24785)

Hi,
*Note: use the table to get the info needed:
.
Probability baby is female from data given = 51/116
.
Probability baby has black hair from data given = ((28 + 16)/116) = 44/116
P(female or has black hair) = 51/116 + 44/116